OpenStudy (anonymous):

When solving x^2 – 12x = –13 by completing the square, what value is added to both sides of the equation? Answer A. 6 B. 144 C. 36 D. –6 When solving x^2 – 12x = –13 by completing the square, what value is added to both sides of the equation? Answer A. 6 B. 144 C. 36 D. –6 @Mathematics

OpenStudy (hoblos):

36

OpenStudy (anonymous):

C

OpenStudy (anonymous):

Thank you! (:

OpenStudy (hoblos):

:D

hero (hero):

(b/2)^2 = (12/2)^2 = (6)^2 = 36